OpenStudy (anonymous):

tan 21 degrees =9/x some please help !!!!!!!!!!! round to nearest tenth

OpenStudy (austinl):

\(\large{\tan(21)=\dfrac{9}{x}}\) \(\large{x=\dfrac{9}{\tan(21)}}\) Plug that into a calulator and make sure you are in degree mode.

OpenStudy (anonymous):

so how would i put it into the calculator

OpenStudy (austinl):

\(\large{9\div \tan(21)}\)

OpenStudy (anonymous):

i got 23.4

OpenStudy (austinl):

Correct!

OpenStudy (anonymous):

another question tan x =0.3

OpenStudy (anonymous):

thank you for your help .. if we need anymore help i will sure come to you

OpenStudy (austinl):

\(\large{\tan(x)=0.3}\) \(\large{x=\tan^{-1}(0.3)}\)

OpenStudy (anonymous):

if i need anymore help can i come to you ??

OpenStudy (austinl):

Yeah, sure. Just tag me in the question like this, @imani1993

OpenStudy (anonymous):

cos 39 degrees =21/m @austinL

OpenStudy (austinl):

\(\large{\cos(39)=\dfrac{21}{m}}\) \(\large{m=\dfrac{21}{\cos(39)}}\)

OpenStudy (anonymous):